what is the center of 4x^2+4y^2+32y=0
Answer by lwsshak3(11628)

what is the center of 4x^2+4y^2+32y=0
..
4x^2+4y^2+32y=0
divide by 4
x^2+y^2+8y=0
complete the square
x^2+(y^2+8+16)-16=0
x^2+(y+4)^2=16
This is a circle with center at (0,-4) and radius of 4